# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v2

@@ Abstract
We study the problem of robot navigation in dense and interactive crowds with environmental constraints such as corridors and furniture. Previous methods fail to consider all types of interactions among agents and obstacles, leading to unsafe and inefficient robot paths. In this article, we leverage a graph-based representation of crowded and constrained scenarios and propose a structured framework to learn robot navigation policies with deep reinforcement learning. We first split the representations of different components in the environment, and propose a heterogeneous spatio-temporal (st) graph to model distinct interactions among humans, robots, and obstacles. Based on the heterogeneous st-graph, we propose HEIGHT, a novel navigation policy network architecture with different components to capture heterogeneous interactions among entities through space and time. HEIGHT utilizes attention mechanisms to prioritize important interactions and a recurrent network to track changes in the dynamic scene over time, encouraging the robot to avoid collisions adaptively. Through extensive simulation and real-world experiments, we demonstrate that HEIGHT outperforms state-of-the-art baselines in terms of success and efficiency in challenging navigation scenarios. Furthermore, in simulation, we demonstrate that our pipeline achieves better zero-shot generalization capability than previous works when the densities of humans and obstacles change. More information is available at https://sites.google.com/view/crowdnav-height/home.
Figure 1: A heterogeneous graph aids the robot’s spatio-temporal reasoning in a crowded and constrained environment. The colored arrows denote robot-human (RH), human-human (HH), and obstacle-agent (OA) interactions. The opaque arrows are the more important interactions while the transparent arrows are the less important ones. The white arrows indicate the front direction of the robot. At each timestep $t$, the robot reasons about these interactions, focuses on the important ones, and makes decisions.
@@ I Introduction
Rising to these challenges, previous works have explored various approaches for robot crowd navigation in constrained spaces [1, 2, 3]. However, these works typically have one of two limitations: (1) They assume agents move in an open space without obstacles [4, 3, 5]; (2) Their scene representations or navigation algorithms do not differentiate between dynamic and static obstacles, and thus the robot has difficulties taking adaptive strategies to avoid collisions [1, 6, 7, 8]. To address these limitations, we ask the following research question: How can a robot navigation policy represent and reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions?
To answer this question, we propose a framework that leverages the heterogeneity of interactions in crowded and constrained scenarios in both scene representation and navigation policy design. First, we use light weight perception modules to extract separate human and obstacle representations that are essential for navigation. The split representations are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, inspired by recent breakthroughs in spatio-temporal (st) networks for crowd navigation [8, 9, 10, 11], we decompose the scenario into a heterogeneous st-graph with different types of edges to represent different types of interactions among the robot, detected humans, and observed obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases.
@@
    We propose a heterogeneous spatio-temporal graph (st-graph) to effectively model the pairwise interactions among all agents and entities in crowded and constrained scenarios.
@@
    Simulation experiments with dense crowds and dense obstacles demonstrate that our method outperforms previous state-of-the-art methods in unseen obstacle layouts. In addition, our method demonstrates better generalization to out-of-distribution environments with different human and obstacle densities.
@@
    We successfully transfer the robot policy learned in a low-fidelity simulator to challenging real-world, everyday dynamic environments.
@@ II Related Works
In this section, we first review the literature of robot crowd navigation, which is divided into model-based and learning-based approaches. Then, we discuss previous crowd navigation efforts in constrained spaces. Finally, we review graph attention mechanism with a focus on the usage of attention networks in multi-agent interaction modeling.
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [1, 12, 13, 14]. Model-based approaches have explored various mathematical models to optimize robot actions [15, 2, 1, 16, 17, 18]. As an early example, ROS navigation stack [19] uses a cost map for global planning and dynamic window approach (DWA) for local planning [1]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [20]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [15, 2]. SF models the interactions between the robot and other agents using attractive and repulsive forces [16]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [21, 20]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [22, 23]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Learning-based approaches have been widely used for navigation in dynamic environments to reduce hyperparameter tuning efforts and the number of assumptions. One example is supervised learning from expert demonstrations of desired behaviors, where the expert ranges from model-based policies [24, 25, 26], human teleoperators in simulators [27], to real pedestrians [28, 29]. Supervised learning does not require explorations of the state and action spaces of the robot, yet the performance of learned policy is limited by the quality of expert demonstrations.
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [30]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [4, 31, 3, 32, 5]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [8]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [8, 33, 23, 5]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [34]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among entities in the scene, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
@@ II-C Scene representation for constrained environments
To deal with both static constraints and humans, some methods such as DWA [1] and DS-RNN [8, 35] use groups of small circles to represent the contours of both static obstacles and humans. While straightforward, groups of circle representation has the following problems: 1. Humans and obstacles are treated in the same way but have different effects on robot decision making. This design choice leads to suboptimal policy as we will show in Sec. I and VI. 2. Since the length of input vectors is proportional to the number of obstacles and humans, the hyperparameters of navigation algorithms easily overfit to a specific environment setting, causing performance degradation when the densities of obstacles and humans change.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [7, 36, 27, 37, 34]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios requires high-fidelity simulators [38, 39, 40] and/or datasets [41, 28, 42]. Since accurate pedestrian simulation is technically difficult and expensive [40], learning a deployable e2e policy in dense and interactive crowds remains an open challenge. To circumvent this problem, other works leverage processed state information such as occupancy maps [30, 43]. However, compared to human circles or point clouds, occupancy maps expand the state space, which reduces training efficiency. In addition, the discretization of agent positions limits the ability of occupancy maps to capture slow agent movements. For this reason, we develop a structured input representation, consisting of human circles and processed obstacle point cloud, that only includes the most essential information for navigation. This input representation leads to strong performance in both simulation and real-world while only requiring a low-fidelity simulator.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [44, 45, 46]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [44]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [45]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [46, 47, 3, 8, 48]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than encoding the states of all agents with a concatenation operation or an recurrent network [4, 31, 35].
@@
More specifically, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [3, 49, 8, 50]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, another line of works use a homogenous graph network to include both RH and HH interactions [6, 5, 51]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [11] and our experiments in Sec. I. To deal with this issue, some recent works use multiple attention networks to encode RH and HH features separately, leading to improved navigation performance in dense and interactive crowds [10, 9, 5]. However, these works only deal with open-world social navigation and thus ignore the interactions between agents and obstacles, which are common in real-world navigation scenarios.
@@
To handle obstacles, Liu et al. [52] and Chen et al. [11] treat the humans, the robot, and static objects as nodes in graph neural networks to model the spatio-temporal interactions among all observed entities. In these works, representing each obstacle as a geometric shape requires further map processing, such as obstacle segmentation. In contrast, we treat the point cloud consisting of all obstacles as a single node in our st-graph, which reduces both the processing requirement and the size of the graph.
@@
Figure 2: An overview of our pipeline in simulation and real-world. (a) At each timestep $t$ in training and testing, the simulator provides a reward $r^{t}$ and the following observations of the environment: obstacle point cloud $o^{t}$, the robot state $w^{t}$, and the human states $h_{1}^{t},...,h_{n}^{t}$, and masks $M^{t}$ (Sec. IV-B1). These observations serve as inputs to HEIGHT, which outputs a robot action $a^{t}$ that maximizes the future expected return $R^{t}$. The simulator executes the actions of all agents and the loop continues. (b) The testing loop in the real-world is similar to the simulator except the perception modules for obtaning the observations are different and the reward is absent.
@@ III Preliminaries
In this section, we formulate constrained crowd navigation as a Markov Decision Process (MDP) and introduce the scene representation and the reward function.
@@ III-A Problem formulation
We model the constrained crowd navigation scenario as a MDP, defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,\mathcal{S}_{0}\rangle$. Let $w^{t}$ be the robot state which consists of the robot’s position $(p_{x},p_{y})$, velocity $(u_{x},u_{y})$, goal position $(g_{x},g_{y})$, and heading angle $\theta$. Let ${h}_{i}^{t}$ be the current state of the $i$-th detected human at time $t$, which consists of the human’s position and velocity $(p_{x}^{i},p_{y}^{i},u_{x}^{i},u_{y}^{i})$. Let ${o}^{t}$ be the current observation of the static obstacles and walls, which is represented as a 2D point cloud. We define the state $s^{t}\in\mathcal{S}$ of the MDP as $s^{t}=[{w}^{t},{o}^{t},{h}^{t}_{1},...,{h}^{t}_{n}]$ if a total number of $n$ humans are detected at the timestep $t$, where $n$ may change within a range in different timesteps. The IDs of humans are untracked, and their states are ordered by an increasing $L_{2}$ distance to the robot. The state space $\mathcal{S}$ is continuous and is expanded in Sec. III-B.
In each episode, the robot begins at an initial state $s^{0}\in\mathcal{S}_{0}$. As shown in Fig. 2(a), according to its policy $\pi(a^{t}|s^{t})$, the robot takes an action $a^{t}\in\mathcal{A}$ at each timestep $t$. The action of the robot consists of the desired translational and rotational accelerations $a^{t}=[a_{trans}^{t},a_{rot}^{t}]$. The action space $\mathcal{A}$ is discrete: the translational acceleration $a_{trans}\in\{-0.05\,m/s^{2},0\,m/s^{2},0.05\,m/s^{2}\}$ and the rotational acceleration $a_{rot}\in\{-0.1\,rad/s^{2},0\,rad/s^{2},0.1\,rad/s^{2}\}$. The translational and rotational velocity is clipped within $[-0.5\,m/s,0.5\,m/s]$ and $[-1\,rad/s,1\,rad/s]$ respectively. The robot motion is governed by the dynamics of TurtleBot 2i. In return, the robot receives a reward $r^{t}$ (see Sec. III-C for details) and transits to the next state $s^{t+1}$ according to an unknown state transition $\mathcal{P}(\cdot|s^{t},a^{t})$. Meanwhile, all other humans also take actions according to their policies and move to the next states with unknown state transition probabilities. The process continues until the robot reaches its goal, $t$ exceeds the maximum episode length $T=491$ steps, or the robot collides with any humans or static obstacles.

### v3

@@ Abstract
We study the problem of robot navigation in dense and interactive crowds with static constraints such as corridors and furniture. Previous methods fail to consider all types of spatial and temporal interactions among agents and obstacles, leading to unsafe and inefficient robot paths. In this article, we leverage a graph-based representation of crowded and constrained scenarios and propose a structured framework to learn robot navigation policies with deep reinforcement learning. We first split the representations of different inputs, and propose a heterogeneous spatio-temporal (st) graph to model distinct interactions among humans, robots, and obstacles. Based on the heterogeneous st-graph, we propose HEIGHT, a novel navigation policy network architecture with different components to capture heterogeneous interactions through space and time. HEIGHT utilizes attention mechanisms to prioritize important interactions and a recurrent network to track changes in the dynamic scene over time, encouraging the robot to avoid collisions adaptively. Through extensive simulation and real-world experiments, we demonstrate that HEIGHT outperforms state-of-the-art baselines in terms of success, navigation time, and generalization to domain shifts in challenging navigation scenarios. More information is available at https://sites.google.com/view/crowdnav-height/home.
Figure 1: A heterogeneous graph aids the robot’s spatio-temporal reasoning in a crowded and constrained environment. The colored arrows denote robot-human (RH), human-human ($H$), and obstacle-agent (OA) interactions. The opaque arrows are the more important interactions, while the transparent arrows are the less important ones. The white arrows indicate the front direction of the robot and its length indicates the robot speed. At each timestep $t$, the robot reasons about these interactions, focuses on the important ones, and makes decisions.
footnotetext: Manuscript received July, 10, 2023; Revised August 8, 2025 and October 12, 2025; Accepted November 13, 2025; Date of publication December 15, 2025; Date of current version December 15, 2025. This article was recommended for publication by Associate Editor and Editor Chenguang Yang upon evaluation of the reviewers’ comments. (Corresponding author: Shuijing Liu.)
\@IEEEabskeysecsize
Note to Practitioners
\@IEEEabskeysecsize\@IEEEgobbleleadPARNLSP
This work is motivated by the challenge of safe and efficient robot navigation in indoor environments with people, such as homes and offices. Conventional navigation methods make simplistic assumptions about interactions among agents and entities, which leads to unsafe or unnatural robot behavior in dynamic settings. Previous machine learning methods either ignore or use ineffective approaches to handle static obstacles—limiting their real-world applicability. For a practical system to reason about both dynamic people and static constraints simultaneously, the following two factors are critical: how the robot represents its surroundings and how it make decisions. To represent the dynamic environment, our method treats people and obstacles differently and keeps the most essential information as inputs. On the architectural side, we propose a structured graph-based network, which consists of specialized models to explicitly models different types of interactions among all observed entities. In challenging simulation benchmarks, our approach outperforms alternative designs that are commonly used in existing works. We also demonstrate strong real-world deployment in everyday environments. To improve deployment, a future direction is continual learning in the real-world.
@@ I Introduction
While a large body of research on robot crowd navigation studies open-world navigation and ignores static obstacles [chen2019crowd, liu2023intention, wang2023navistar], some previous works have explored various approaches for crowd navigation in constrained spaces [fox1997dynamic, van2008reciprocal, perez2021robot]. However, these works typically have one of two limitations: (1) They do not differentiate between dynamic and static obstacles, and thus the robot has difficulties taking adaptive strategies to avoid collisions [fox1997dynamic, chen2019relational, perez2021robot, liu2020decentralized]; (2) The representations of scene, especially static obstacles, are inefficient or redundant as inputs to the navigation algorithm. To address these limitations, we ask the following research question: How can a robot navigation policy represent and reason about diverse interactions in crowded and constrained environments to adaptively avoid collisions?
To answer this question, we propose a framework that leverages the heterogeneity of interactions in crowded and constrained scenarios in both scene representation and navigation policy design. First, we use light weight perception modules to extract separate human and obstacle representations that are essential for navigation, such as relative positions w.r.t. the robot. The split representations are processed and fed separately into the reinforcement learning (RL)-based navigation pipeline. Then, inspired by recent breakthroughs in spatio-temporal (st) networks for crowd navigation [liu2020decentralized, wang2023navistar, liu2023intention, chen2024socially], we decompose the scenario into a heterogeneous st-graph with different types of edges to represent different types of interactions among the robot, detected humans, and observed obstacles, as shown in the colored arrows in Fig. 1. Finally, we convert the heterogeneous st-graph into a HEterogeneous Interaction GrapH Transformer (HEIGHT), a robot policy network consisting of different modules to parameterize the various spatio-temporal interactions. Specifically, we use two separate attention networks to address the different effects of robot-human (RH) and human-human (HH) interactions. The attention networks enable the robot to pay more attention to the important interactions, leading to a low collision-rate even as the number of humans increases.
@@
    We propose a heterogeneous spatio-temporal (st) graph to effectively model the pairwise interactions among all agents and entities in crowded and constrained scenarios.
@@
    In simulation benchmarks with dense crowds and dense obstacles, our method outperforms previous state-of-the-art methods in unseen obstacle layouts. In addition, our method demonstrates better generalization w.r.t. different human and obstacle densities.
@@
    We successfully transfer the robot policy learned in a low-fidelity simulator to challenging real-world, everyday dynamic environments such as narrow corridors and densely packed public spaces.
@@ II Related Works
In this section, we first review the literature of robot crowd navigation (Sec. II-A and II-B). Then, we narrow down to crowd navigation efforts in constrained spaces (Sec. II-C). Finally, we review graph attention mechanism with a focus on its usage in multi-agent interaction modeling (Sec. II-D).
@@ II-A Model-based methods
Robot navigation in human crowds is particularly challenging and has been studied for decades [fox1997dynamic, hoy2015algorithms, savkin2014seeking, mavrogiannis2021core]. Model-based approaches have explored various mathematical models to optimize robot actions [van2011reciprocal, van2008reciprocal, fox1997dynamic, helbing1995social, huber2022avoiding, mavrogiannis2023winding]. As an early example, ROS navigation stack [ros_navigation_stack] uses a cost map for global planning and dynamic window approach (DWA) for local planning [fox1997dynamic]. DWA searches for the optimal velocity that brings the robot to the goal with the maximum clearance from any obstacle while obeying the dynamics of the robot. By treating humans as obstacles, DWA exhibits myopic behaviors such as oscillatory paths in dynamic environments [dobrevski2024dynamic]. As a step forward, optimal reciprocal collision avoidance (ORCA) and social force account for the velocities of agents. ORCA models other agents as velocity obstacles and assumes that agents avoid each other under the reciprocal rule [van2011reciprocal, van2008reciprocal]. SF models the interactions between the robot and other agents using attractive and repulsive forces [helbing1995social]. However, the hyperparameters of the model-based approaches are sensitive to crowd behaviors and thus need to be tuned carefully to ensure good performance [long2018towards, dobrevski2024dynamic]. In addition, model-based methods are prone to failures, such as the freezing problem, if the assumptions such as the reciprocal rule are broken [trautman2010unfreezing, liu2024relaxing]. In contrast, while our method also models these interactions, we learn the hyperparameters of the model from trial and error with minimal assumptions on human behaviors posed by model-based methods.
@@ II-B Learning-based methods
Learning-based approaches have been widely used for navigation in dynamic environments to reduce hyperparameter tuning efforts and the number of assumptions. One example is supervised learning from expert demonstrations of desired behaviors, where the expert ranges from model-based policies [long2017deep, tai2018socially, xie2021towards], human teleoperators in simulators [pokle2019deep], to real pedestrians [karnan2022scand, chandra2024towards]. Supervised learning does not require explorations of the state and action spaces of the robot, yet the performance of learned policy is limited by the quality of expert demonstrations.
Another line of work takes advantage of crowd simulators and learns policies with RL. Through trial and error, RL has the potential to learn robot behaviors that outperform model-based approaches and expert demonstrations [xie2023drlvo]. For example, Deep V-Learning first uses supervised learning with ORCA as the expert and then uses RL to learn a value function for path planning [chen2017decentralized, everett2018motion, chen2019crowd, chen2020robot_gaze, yang2023st2]. However, Deep V-Learning assumes that state transitions of all agents are known without uncertainty. In addition, since the networks are pre-trained with supervised learning, they share the same disadvantages with OCRA, which are hard to correct by RL [liu2020decentralized]. To address these problems, more recent efforts leverage model-free RL without supervised learning or assumptions on state transitions [liu2020decentralized, mun2023occlusion, liu2024relaxing, yang2023st2]. Since the state transition probability of humans is uncertain, directly learning a policy network without explicitly modeling state transitions is more suitable for navigation [sathyamoorthy2020densecavoid]. However, the aforementioned RL works have at least one of the following two problems: (1) They focus on navigation in open spaces and isolate agents from static constraints, posing difficulties when deploying robots in the real-world; (2) They ignore all or part of interactions among entities in the scene, which are important for robot navigation in dense crowds and highly constrained environments. We discuss how prior works address the above two problems in Sec. II-C and Sec. II-D respectively.
@@ II-C Scene representation for constrained environments
To deal with both static constraints and humans, some methods such as DWA [fox1997dynamic] and DS-RNN [liu2020decentralized, chen2024decentralized] use groups of small circles to represent the contours of both static obstacles and humans. While straightforward, groups of circle representation has the following problems: 1. Humans and obstacles are treated in the same way but have different effects on robot decision making. This design choice leads to suboptimal policy as we will show in Sec. V and VI. 2. Since the length of input vectors is proportional to the number of obstacles and humans, the hyperparameters of navigation algorithms easily overfit to a specific environment setting, causing performance degradation when the densities of obstacles and humans change.
Other learning-based approaches use raw sensor images or point clouds to represent the whole environment [perez2021robot, dugas2022navdreams, pokle2019deep, zheng2022hierarchical, sathyamoorthy2020densecavoid]. These end-to-end (e2e) pipelines have made promising progress in simulation. However, generalizing these e2e methods to real-world scenarios requires high-fidelity simulators [tsoi2022sean2, puig2024habitat, Kstner2024Arena3A] and/or datasets [martin2021jrdb, karnan2022scand, hirose2023sacson]. Since accurate pedestrian simulation is technically difficult and expensive [Kstner2024Arena3A], learning a deployable e2e policy in dense and interactive crowds remains an open challenge. To circumvent this problem, other works leverage processed state information such as occupancy maps [xie2023drlvo, yang2023rmrl]. However, compared to human circles or point clouds, occupancy maps expand the state space, which reduces training efficiency. In addition, the discretization of agent positions limits the ability of occupancy maps to capture slow agent movements. For this reason, we develop a structured input representation, consisting of human circles and processed obstacle point cloud, that only includes the most essential information for navigation. This input representation leads to strong performance in both simulation and real-world while only requiring a low-fidelity simulator.
@@ II-D Graph attention networks for multi-agent interactions
In recent years, attention mechanisms have demonstrated success in various fields [vaswani2017attention, velickovic2018graph, vemula2018social]. Vaswani et al. propose a transformer with self-attention mechanism that achieves state-of-the-art performance in machine translation [vaswani2017attention]. Later, graph attention networks show the effectiveness of attention on learning relationships of data with graphical structures such as citation networks [velickovic2018graph]. Inspired by these works, researchers in trajectory prediction and crowd navigation have found that attention networks are also well-suited to capture interactions amongst agents and entities, which contain essential information for multi-agent tasks [vemula2018social, huang2019stgat, chen2019crowd, liu2020decentralized, hasan2022meta]. For each agent, these works compute attention scores for all neighboring agents. Due to the permutation invariance property, attention scores better capture the importance of pair-wise relationships than encoding the states of all agents with a concatenation operation or an recurrent network [chen2017decentralized, everett2018motion, chen2024decentralized].
@@
More specifically, a line of works uses a robot-human attention network to determine the relative importance of each human to the robot [chen2019crowd, leurent2019social, liu2020decentralized, liu2021learning]. However, interactions among humans, which can also influence the robot, are not explicitly modeled. To this end, another line of works use a homogenous graph network to include both RH and HH interactions [chen2019relational, yang2023st2, liu2024sample]. However, since these works feed RH and HH features to a single attention network, the resulting robot policy has difficulty reasoning the specificity of each type of feature, which limits the robot’s ability to adapt to different interactions, as demonstrated in [chen2024socially] and our experiments in Sec. V. To deal with this issue, some recent works use multiple attention networks to encode RH and HH features separately, leading to improved navigation performance in dense and interactive crowds [liu2023intention, wang2023navistar, yang2023st2]. However, these works only deal with open-world social navigation and thus ignore the interactions between agents and obstacles, which are common in real-world navigation scenarios.
@@
To handle obstacles, Liu et al. [liu2023graph] and Chen et al. [chen2024socially] treat the humans, the robot, and static objects as nodes in graph neural networks to model the spatio-temporal interactions among all observed entities. Different from Chen et al. [chen2024socially] where a robot needs to reach an object such as a chair, the robot in this work only needs to avoid all objects. Thus, Chen et al. [chen2024socially] represent static obstacles as a semantic map with a class label for each object, whereas we represent all observed static obstacles as a point cloud. In Liu et al. [liu2023graph], representing each obstacle as a geometric shape requires further map processing, such as obstacle segmentation. Thus, we treat the point cloud consisting of all obstacles as a single node in our st-graph, which reduces both the processing requirement and the size of the graph.
@@
Figure 2: An overview of our pipeline in simulation and real-world. (a) At each timestep $t$ in training and testing, the simulator provides a reward $r^{t}$ and the following observations of the environment: obstacle point cloud $o^{t}$, the robot state $w^{t}$, and the human states $h_{1}^{t},...,h_{n}^{t}$, and masks $M^{t}$ (Sec. IV-B1). These observations serve as inputs to HEIGHT, which outputs a robot action $a^{t}$ that maximizes the future expected return $R^{t}$. The simulator executes the actions of all agents and the loop continues. (b) The testing loop in the real-world is similar to the simulator except the perception modules for obtaining the observations are different and the reward is absent.
@@ III Preliminaries
In this section, we present our problem formulation, scene representation, and reward function.
@@ III-A Problem formulation
We model the constrained crowd navigation scenario as a Markov Decision Process (MDP), defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{P},R,\gamma,\mathcal{S}_{0}\rangle$. Let $w^{t}$ be the robot state which consists of the robot’s position $(p_{x},p_{y})$, instantaneous velocity $(u_{x},u_{y})$, goal position $(g_{x},g_{y})$, and heading angle $\theta$ in world frame. Let ${h}_{i}^{t}$ be the current state of the $i$-th detected human at time $t$, which consists of the human’s relative position w.r.t. the robot and its instantaneous velocity $(p_{x}^{i}-p_{x},p_{y}^{i}-p_{y},u_{x}^{i},u_{y}^{i})$ in world frame in simulation experiments. In real world experiments, the state of the $i$-th human only includes its relative position $h^{t}_{i}=(p_{x}^{i}-p_{x},p_{y}^{i}-p_{y})$ since accurate estimation of human velocities are difficult to obtain. Let ${o}^{t}$ be the current observation of the static obstacles and walls, which is represented as a 2D point cloud. We define the state $s^{t}\in\mathcal{S}$ of the MDP as $s^{t}=[{w}^{t},{o}^{t},{h}^{t}_{1},...,{h}^{t}_{n}]$ if a total number of $n$ humans are detected at the timestep $t$, where $n$ may change within a range in different timesteps. The IDs of humans are untracked, and their states are ordered by an increasing $L_{2}$ distance to the robot. We add a small Gaussian noise to the ground truth positions and velocities of all agents to compensate for perception errors. The state space $\mathcal{S}$ is continuous and is expanded in Sec. III-B.
In each episode, the robot begins at an initial state $s^{0}\in\mathcal{S}_{0}$. As shown in Fig. 2(a), according to its policy $\pi(a^{t}|s^{t})$, the robot takes an action $a^{t}\in\mathcal{A}$ at each timestep $t$. The action of the robot consists of the desired translational and rotational accelerations $a^{t}=[a_{trans}^{t},a_{rot}^{t}]$. The action space $\mathcal{A}$ is discrete: the translational acceleration $a_{trans}\in\{-0.05\,m/s^{2},0\,m/s^{2},0.05\,m/s^{2}\}$ and the rotational acceleration $a_{rot}\in\{-0.1\,rad/s^{2},0\,rad/s^{2},0.1\,rad/s^{2}\}$. The robot translational velocity is $v_{trans}^{t}=v_{trans}^{t-1}+a_{trans}^{t}\times\Delta t\in[-0.5\,m/s,0.5\,m/s]$ and rotational velocity is $v_{rot}^{t}=v_{rot}^{t-1}+a_{rot}^{t}\times\Delta t\in[-1\,rad/s,1\,rad/s]$. The robot motion is governed by differential drive dynamics. In return, the robot receives a reward $r^{t}$ (see Sec. III-C for details) and transits to the next state $s^{t+1}$ according to an unknown state transition $\mathcal{P}(\cdot|s^{t},a^{t})$. Meanwhile, all other humans also take actions according to their policies and move to the next states with unknown state transition probabilities. The process continues until the robot reaches its goal, $t$ exceeds the maximum episode length $T=491$ steps, or the robot collides with any humans or static obstacles.